Magazine issued $300,000 of​ 15-year, 6% callable bonds payable on July​ 31, 2018​, at 93. On July​ 31, 2021​, Parkview called the bonds at 101. Assume annual interest payments.

Requirements

1.

Without making journal​ entries, compute the carrying amount of the bonds payable at July​ 31, 2021.

2.

Assume all amortization has been recorded properly. Journalize the retirement of the bonds on July​ 31, 2021. No explanation is required
Answer #1

1)Issue price of Bond =Par value * issue price per 100

                    = 300000*93/100

                    = 279000

Discount on bond issued = 300000- 279000= 21000

Amortization of bond discount per year = 21000/15 years = 1400 per year

Total amortization till 31July 2021 (31July 2018-31July 2021 = 3years) = 1400*3=4200

Carrying value of Bond as on 1July 2021 =Issue price +Discount amortized till date

                                    = 279000 + 4200

                                    = 283200

2)

Discount on bond (un-amortized) =21000-4200 = 16800

Loss on retirement of bond =price paid -carrying value

                          = 303000 - 283200

                         = 19800

DATE ACCOUNT TITLE DEBIT CREDIT
31 July 2021 Bond payable 300000
Loss on retirement of bond 19800
Discount on bond payable 16800
Cash (300000*101/100) 303000